# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ 6. $RQ_{2}$: Antecedents of Trust
Reliability ”My trust is based on the tool’s ability to consistently and reliably generate correct, understandable outputs.” (Johnson et al., 2023; Lo, 2023; Liu et al., 2024; Maninger et al., 2024)
@@
Transparency Normative Metrics ”Trust in LLMs could be interpreted as the percentage of accurate code generated within an extended time frame of use.” (Key et al., 2023; Spiess et al., 2024; Lo, 2023; Widder et al., 2021b; Liu et al., 2024)
XAI Features ”I trust the tools if I can understand the snippet and how it functions for myself.” (Widder et al., 2021b; Johnson et al., 2023; Key et al., 2023; Weisz et al., 2021; Noller et al., 2022)
Interaction Customization ”It’s difficult to capture the requirements through a simple prompt.” (Perry et al., 2023; Johnson et al., 2023; Widder et al., 2021b)
@@ 6.1. Phase1: SLR Results
From the literature, we extracted a large collection of antecedents that impact human trust in using LLM4SE. We categorize these antecedents into two main groups: model-specific trustworthiness attributes, which are inherent to the model, and user-centric trust attributes, which are extrinsic to the model. Table 2 provides a comprehensive overview of the taxonomy we developed, constructed using the MATCH model  (Liao and Sundar, 2022) as a guiding framework. 61% (n=11 out of 18) of the studies focused on antecedents inherent to the model during the design phase; Conversely, 22% (n=4 out of 18) explored how attributes associated with interface interaction help deliver model trustworthiness to stakeholders; 17% (n=3 out of 18) explored both aspects.
@@ 6.1.1. Model-Specific Attribute
  • Ability: Ability is an overarching term for LLMs’ performance or competence (Wang et al., 2024c). We identified the three most popular performance-related antecedents from our literature: accuracy, efficiency, and reliability. Accuracy is the degree to which the output matches the expected results in terms of correctness and precision (Key et al., 2023; Lo, 2023; Johnson et al., 2023). Efficiency refers to the software’s capability to use computational resources judiciously and effectively (Lo, 2023). Reliability is the ability of the software to function correctly and consistently under various conditions for an extended period (Johnson et al., 2023; Lo, 2023; Liu et al., 2024). Given the probabilistic nature of LLMs, models occasionally generate incorrect or insecure outputs. This inherent unreliability poses a key challenge to establishing trust in these models (Maninger et al., 2024). Reliability also includes the model’s ability to handle unexpected input gracefully and have appropriate fault tolerance (Liu et al., 2024).
@@
  • Benevolence: A trustworthy model should prioritize acting in the best interest of users by aligning its goals with theirs (Wang et al., 2024c). This alignment can be categorized into two levels: short-term goal alignment and long-term goal alignment. In the short term, models must effectively address users’ immediate needs, which can vary among individuals in different situations. These needs may include offering clear advantages in usage (Johnson et al., 2023), improving productivity during the coding process (Wang et al., 2024c), and meeting specific functional expectations (Johnson et al., 2023). While achieving short-term goals is crucial for fostering initial trust, it alone is insufficient to ensure lasting trustworthiness.
@@
    Beyond immediate outputs, models must align with the long-term goals of users. For example, developers often rely on LLMs to quickly and accurately achieve their short-term objectives. However, they are also concerned about maintaining and improving their programming skills over time, as excessive dependence on LLMs may lead to skill degradation  (Wang et al., 2024c). To address this concern, LLMs should incorporate educational content into their generated outputs, fostering learning and supporting developers in strengthening their expertise (Johnson et al., 2023). However, current models lack mechanisms that allow developers to clearly specify both their short-term and long-term goals, leading to a misalignment between the developer’s expectations and the outputs generated by the AI (Wang et al., 2024c).
@@
  • Integrity: This attribute pertains to the training process of LLMs, including the development practice was safe and secure, under ethical guidelines (Maninger et al., 2024; Lo, 2023; Johnson et al., 2023). One study highlights the critical link between training data quality and model trustworthiness, emphasizing how issues such as biases, outdated information, or security vulnerabilities can propagate into the model’s outputs, directly impacting its integrity (Wang et al., 2024a). Regarding the training data, models are expected to reduce bias, which can be achieved by updating training data for new tasks and curating real-world datasets (Russo, 2024; Lo, 2023; Maninger et al., 2024).
@@ 6.1.2. User-Centric Attributes
User-Centric Trust Attributes are trust factors that arise from a user’s interaction with the system, including transparency, explainability, usability, and the ability to align the results with user requirements and expectations. In an ideal scenario, model-specific attributes are effectively communicated to users. However, in typical usage, users cannot directly assess a model’s trustworthiness; instead, they infer it through various affordances. Designing a trustworthy model involves creating trust affordances that deliver the actual capability of the model, so users can form appropriate trust without overestimating or underestimating the model’s capability. In this section we, we present results related to user-centric attributes.
@@
  • Transparency: Transparency is an affordance employed by developers to signal to stakeholders that their trust in the system is well-founded (Kästner et al., 2021). By offering clear justifications, transparency helps users believe that the system operates as intended. In our analysis, normative metrics and XAI features are the most emphasized factors for achieving transparency.
@@
    Efficiency is another key aspect evaluated through normative metrics. Efficiency metrics assess the time taken by the model to execute specific tasks, providing users with clear expectations about response times and helping to prevent surprises. This transparency about response time also builds confidence in the model’s practical usability. Similarly, reliability is often measured by the mean time between failures (MTBF), assuring users that the model consistently delivers accurate outputs without frequent errors (Widder et al., 2021b).
    Normative metrics are closely tied to transparency because they offer users the data they need to gauge the model’s performance. A well-calibrated model, for instance, provides probabilistic estimates that align closely with actual outcomes, enhancing trust by showing that the model’s predictions are accurate, effective, and reliable  (Key et al., 2023; Spiess et al., 2024). However, the current models suffer from issues like data duplication, leading to over-optimistic results  (Liu et al., 2024).
    Alongside normative metrics, XAI features serve as another essential mechanism for enhancing transparency. XAI enhances transparency by offering insights into how and why decisions are made, helping users understand the model’s reasoning and inner workings. Relevant documentation and training materials that clarify the rationale behind the model’s outputs can further aid in building trust  (Widder et al., 2021b). Studies have shown that providing clear comments and specifications in code helps users understand a program’s behavior  (Johnson et al., 2023).
@@
    When users engage directly with the model, they often assess its customizability—the extent to which the model adapts to user preferences and feedback. Studies show that users value models that can evolve based on their input, reinforcing perceptions of the model’s ability and responsiveness  (Johnson et al., 2023; Widder et al., 2021b). Despite these affordances, limitations in prompt comprehension remain a significant barrier. Users who can effectively articulate and refine their prompts tend to achieve better outcomes, while novice users often struggle with the learning curve of prompt engineering  (Perry et al., 2023).
    Beyond direct engagement, users frequently turn to the broader community to inform their trust judgments. Socialization—interacting with user communities and leveraging collective experiences—provides valuable insights into the model’s strengths and limitations  (Cheng et al., 2023). Observing community feedback allows users to gain a more comprehensive understanding of the model’s capabilities, often supplementing the limited perspective they might form through personal use (Cheng et al., 2023). Positive examples, shared strategies, and troubleshooting tips from others help build trust by demonstrating the model’s practical value across diverse scenarios (Cheng et al., 2023).
@@ 6.2. Phase2: Survey Study Results
In the survey study, we asked the respondents to select one or more antecedents that would influence their trust in LLM4SE tasks. These factors can be categorized into:
@@
[…]
[Trust Antecedents]Trust Antecedents Selection Based on SE Tasks
Figure 2. Trust Antecedents Selection Based on SE Tasks
Fig. 2 is a grouped bar chart that visualizes the selection of trust antecedents across three tasks. In general, we observe that model trustworthiness attributes are valued more than human-model interaction attributes, regardless of the SE tasks. Accuracy is the most selected trust antecedent across all tasks (94% for code generation; 100% for test case generation and program repair). However, preferences for the second most important antecedents vary by task: for code generation, controllability (58.8%) and robustness (58.8%) are equally valued; for test-case generation, controllability (75%) is preferred; and for program repair, interpretability (72.7%) is valued.
@@
[…]
[Trust Antecedents]Trust Antecedents Selection Based on Years of Experience with LLMs
Figure 3. Trust Antecedents Selection Based on Years of Experience with LLMs
Fig. 3 is a heat map showing the relationship between user experiences using LLM4SE and their trust antecedent selection. The percentage is calculated as the weighted average of respondents across three SE tasks.
The visualization shows accuracy and robustness are highly valued across all experience levels. Conversely, community-associated factors (community engagement, endorsement, source reputation) are not valued by new users but gain attention with increased experience, though still less than model-related attributes. This suggests that users base trust more on direct model interaction than external validation.
A clear trend also emerges based on user experience: experienced users place less emphasis on the model’s controllability and ethical compliance. This could imply that with increasing familiarity, users developed a level of trust that mitigates concerns about how to use the tool effectively and the model’s benevolence. On the other hand, experienced users increasingly value the interpretability of the generated output and prefer models that seamlessly integrate into their established procedure.
[…]
[Trust Antecedents]Trust Antecedents based on years of Experience in SE
Figure 4. Trust Antecedents Selection Based on Years of Experience in SE
We observe a similar result from Fig. 4, a heat map between users’ year of expertise in SE and their trust antecedent selection. However, we find that there is no significant relationship between higher levels of expertise in SE and their selection of community-related factors. Interestingly, we observe that intermediates pay significantly less emphasis on robustness compared to novices and experts.
@@ 6.3. Complementary Analysis of Trust Antecedents
Trust in AI systems, including LLMs, is shaped by multiple interrelated factors, spanning technical performance, user perception, and ethical considerations. In AI research, accuracy is widely recognized as a fundamental trust factor across domains  (Khandelwal et al., 2020; Hendrycks et al., 2021; Yin et al., 2023). In SE, ensuring accuracy and reliability is essential to preventing security vulnerabilities  (Hou and Jansen, 2022; Ahuja et al., 2020; Hassan et al., 2024; Baldassarre et al., 2024a; Serban et al., 2021; Rutinowski et al., 2024b). However, trust is not solely dependent on accuracy; overconfidence in the accuracy of the model can hinder trust by making unreliable outputs appear convincing  (Huang et al., 2023). This phenomenon is observed in various AI applications, where users may develop misplaced trust in systems despite their inherent unreliability  (Banovic et al., 2023).
@@
Beyond accuracy, trust is also influenced by the model’s ability to balance technical performance with inclusivity and multilingual capabilities. NLP research highlights that ensuring unbiased and culturally aware outputs remains a challenge when optimizing for high performance across different linguistic contexts  (Cabrero-Daniel and Sanagustín Cabrero, 2023; Zhang et al., 2023; Wang et al., 2024b). This trade-off between precision and inclusivity is particularly relevant for SE, where domain-specific terminology may not generalize well across different programming cultures.
@@
Robustness is another critical factor in trust formation. AI models, including LLMs, often struggle with maintaining consistent performance across similar inputs. Studies show that even minor perturbations in prompts can lead to significantly different outputs  (Ye et al., 2023; Mehrotra et al., 2024), and evaluations such as PromptBench highlight LLMs’ sensitivity to word-level variations  (Zhu et al., 2024). Similar robustness challenges have been identified in NLP tasks, where adversarial manipulations can undermine model reliability  (Wang et al., 2022; Chen and Hsieh, 2023).
@@
Transparency and accountability further shape trust in AI systems. Research suggests that users develop greater trust when they perceive AI as competent, benevolent, and transparent in its decision-making  (Baughan et al., 2023; Jakesch et al., 2019; Zerick et al., 2024; Gambo et al., 2024). However, the black-box nature of LLMs and corporate secrecy hinder transparency  (Wu et al., 2022; Liao and Wortman Vaughan, 2024; Langer et al., 2021). Techniques such as uncertainty expressions (e.g., “I’m not sure”) can improve trust by signaling limitations  (Kim et al., 2024), while XAI approaches offer interpretability at the potential cost of performance  (Wu et al., 2022). The regulatory landscape also plays a role, with frameworks like the EU AI Act emphasizing responsibility  (EUw, [n. d.]; Jobin et al., 2019; Floridi and Cowls, 2019; Nissenbaum, 1996; Stettinger et al., 2024), yet leaving questions of liability for AI errors unresolved  (Nissenbaum, 1996; Lassiter and Fleischmann, 2024). Proposed solutions, such as blockchain-based traceability, could enhance accountability in AI decision-making  (Zhang et al., 2024).
@@
Finally, ethical considerations remain an essential, albeit sometimes secondary, factor in trust formation. While performance and usability often take priority, fairness in AI is critical to long-term trustworthiness. Studies reveal that LLMs can introduce social biases even in code generation tasks  (Liu et al., 2023), and research on models like Codex indicates a trade-off between fairness and performance  (Cui et al., 2024). Trust is also shaped by concerns about AI’s impact on professional development. Similar to educational AI, where trust is linked to fairness and correctness in recommendations  (Amoozadeh et al., 2024), SE users worry that overreliance on LLMs could degrade their programming skills. However, while not a primary focus for experienced users, ethical concerns remain an important aspect of long-term trust in AI  (Baldassarre et al., 2024c).

### v2

@@ 6. $RQ_{2}$: Antecedents of trust
Reliability ”My trust is based on the tool’s ability to consistently and reliably generate correct, understandable outputs.” (Johnson et al., 2023; Lo, 2023; Liu et al., 2024b; Maninger et al., 2024)
@@
Transparency Normative Metrics ”Trust in LLMs could be interpreted as the percentage of accurate code generated within an extended time frame of use.” (Key et al., 2023; Spiess et al., 2024; Lo, 2023; Widder et al., 2021; Liu et al., 2024b)
Explainable AI Features ”I trust the tools if I can understand the snippet and how it functions for myself.” (Widder et al., 2021; Johnson et al., 2023; Key et al., 2023; Weisz et al., 2021; Noller et al., 2022)
Interaction Customization ”It’s difficult to capture the requirements through a simple prompt.” (Perry et al., 2023; Johnson et al., 2023; Widder et al., 2021)
@@ 6.1. Phase1: SLR results
From the literature, we extracted a large collection of antecedents that impact human trust in the use of LLM4SE. We categorize these antecedents into two main groups: model-specific trustworthiness attributes, which are inherent to the model, and user-centric trust attributes, which are extrinsic to the model. Table 3 provides a comprehensive overview of the taxonomy we developed, constructed using the MATCH model (Liao and Sundar, 2022) as a guiding framework. 61% (n=11 out of 18) of the studies focused on the antecedents inherent to the model during the design phase; Conversely, 22% (n=4 out of 18) explored how attributes associated with interface interaction help deliver model trustworthiness to stakeholders; 17% (n=3 out of 18) explored both aspects.
@@ 6.1.1. Model-Specific Attribute
  • Ability: Ability is an overarching term for LLMs’ performance or competence (Wang et al., 2024c). We identified the three most popular performance-related antecedents in our literature: accuracy, efficiency, and reliability. Accuracy is the degree to which the output matches the expected results in terms of accuracy and precision (Key et al., 2023; Lo, 2023; Johnson et al., 2023). Efficiency refers to the ability of the software to use computational resources judiciously and effectively (Lo, 2023). Reliability is the ability of the software to function correctly and consistently under various conditions for an extended period (Johnson et al., 2023; Lo, 2023; Liu et al., 2024b). Given the probabilistic nature of LLMs, models sometimes generate incorrect or insecure outputs. This inherent unreliability poses a key challenge in establishing trust in these models (Maninger et al., 2024). Reliability also includes the model’s ability to handle unexpected input gracefully and to have appropriate fault tolerance (Liu et al., 2024b).
@@
  • Benevolence: A trustworthy model should prioritize acting in the best interest of users by aligning its goals with theirs (Wang et al., 2024c). This alignment can be categorized into two levels: short-term goal alignment and long-term goal alignment. In the short term, models must effectively address users’ immediate needs, which can vary among individuals in different situations. These needs may include offering clear advantages in usage (Johnson et al., 2023), improving productivity during the coding process (Wang et al., 2024c), and meeting specific functional expectations (Johnson et al., 2023). Although achieving short-term goals is crucial for fostering initial trust, it alone is not sufficient to ensure long-lasting trustworthiness.
@@
    Beyond immediate outputs, models must align with the long-term goals of users. For example, developers often rely on LLMs to quickly and accurately achieve their short-term objectives. However, they are also concerned about maintaining and improving their programming skills over time, as excessive dependence on LLMs may lead to skill degradation (Wang et al., 2024c). To address this concern, LLMs should incorporate educational content into their generated outputs, fostering learning and supporting developers in strengthening their expertise (Johnson et al., 2023). However, current models lack mechanisms that allow developers to clearly specify both their short-term and long-term goals, leading to a misalignment between the developer’s expectations and the results generated by AI (Wang et al., 2024c).
@@
  • Integrity: This attribute is related to the LLM training process, including that the development practice was safe and secure, according to ethical guidelines (Maninger et al., 2024; Lo, 2023; Johnson et al., 2023). A study highlights the critical link between training data quality and model trustworthiness, emphasizing how issues such as biases, outdated information, or security vulnerabilities can propagate into model output, directly impacting its integrity (Wang et al., 2024a). Regarding training data, models are expected to reduce bias, which can be achieved by updating training data for new tasks and curating real-world data sets (Russo, 2024; Lo, 2023; Maninger et al., 2024).
@@ 6.1.2. User-Centric Attributes
User-Centric Trust Attributes are trust factors that arise from a user’s interaction with the system, including transparency, explainability, usability and the ability to align the results with user requirements and expectations. In an ideal scenario, model-specific attributes are effectively communicated to users. However, in typical usage, users cannot directly assess a model’s trustworthiness; instead, they infer it through various affordances. Designing a trustworthy model involves creating trust affordances that deliver the actual capability of the model, so users can form appropriate trust without overestimating or underestimating the model’s capability. In this section, we present results related to user-centric attributes.
@@
  • Transparency: Transparency is an affordance employed by developers to signal to stakeholders that their trust in the system is well-founded (Kästner et al., 2021). By offering clear justifications, transparency helps users believe that the system operates as intended. In our analysis, normative metrics and Explainable Artificial Intelligence (XAI) features are the most emphasized factors for achieving transparency.
@@
    Efficiency is another key aspect evaluated through normative metrics. Efficiency metrics assess the time taken by the model to execute specific tasks, providing users with clear expectations about response times and helping to prevent surprises. This transparency about response time also builds confidence in the model’s practical usability. Similarly, reliability is often measured by the Mean Time Between Failures (MTBF), assuring users that the model consistently delivers accurate outputs without frequent errors (Widder et al., 2021).
    Normative metrics are closely tied to transparency because they offer users the data they need to gauge the model’s performance. A well-calibrated model, for instance, provides probabilistic estimates that align closely with actual outcomes, enhancing trust by showing that the model’s predictions are accurate, effective, and reliable (Key et al., 2023; Spiess et al., 2024). However, the current models suffer from issues like data duplication, leading to over-optimistic results (Liu et al., 2024b).
    Alongside normative metrics, XAI features serve as another essential mechanism for enhancing transparency. XAI enhances transparency by offering insights into how and why decisions are made, helping users understand the model’s reasoning and inner workings. Relevant documentation and training materials that clarify the rationale behind the model’s outputs can further aid in building trust (Widder et al., 2021). Studies have shown that providing clear comments and specifications in code helps users understand a program’s behavior (Johnson et al., 2023).
@@
    When users engage directly with the model, they often assess its customizability—the extent to which the model adapts to user preferences and feedback. Studies show that users value models that can evolve based on their input, reinforcing perceptions of the model’s ability and responsiveness (Johnson et al., 2023; Widder et al., 2021). Despite these affordances, limitations in prompt comprehension remain a significant barrier. Users who can effectively articulate and refine their prompts tend to achieve better outcomes, while novice users often struggle with the learning curve of prompt engineering (Perry et al., 2023).
    Beyond direct engagement, users frequently turn to the broader community to inform their trust judgments. Socialization—interacting with user communities and leveraging collective experiences—provides valuable insights into the model’s strengths and limitations (Cheng et al., 2023). Trustworthiness perceptions may also differ based on whether the LLM is open-source or proprietary. As observed by Huang et al. (Huang et al., 2024), proprietary models tend to outperform open-source ones in terms of trustworthiness; however, models like LLaMA2 demonstrate that open-source systems can closely rival their proprietary counterparts. These structural differences in model openness further shape how users interpret and validate feedback from the community. Observing community feedback allows users to gain a more comprehensive understanding of the model’s capabilities, often supplementing the limited perspective they might form through personal use (Cheng et al., 2023). Positive examples, shared strategies, and troubleshooting tips from others help build trust by demonstrating the model’s practical value across diverse scenarios (Cheng et al., 2023).
@@ 6.2. Phase2: Survey Study Results
In the survey study, we asked respondents to select one or more antecedents that would influence their trust in LLM4SE tasks. These factors can be categorized into the following:
@@
Figure 3. Trust Antecedents Selection Based on SE Tasks: Trust Antecedents
Fig. 3 is a grouped bar chart that visualizes the selection of trust antecedents across three tasks. In general, we observe that model trustworthiness attributes are valued more than human-model interaction attributes, regardless of SE tasks. Accuracy is the most selected trust antecedent in all tasks (94% for code generation; 100% for test case generation and program repair). However, preferences for the second most important antecedents vary by task: for code generation, controllability (58.8%) and robustness (58.8%) are equally valued; for test-case generation, controllability (75%) is preferred; and for program repair, interpretability (72.7%) is valued.
@@
Figure 4. Trust Antecedents Selection Based on Years of Experience with LLMs: Trust Antecedents
Fig. 4 is a heat map showing the relationship between user experiences using LLM4SE and their trust antecedent selection. The percentage is calculated as the weighted average of respondents in three SE tasks.
The visualization shows that accuracy and robustness are highly valued across all experience levels. Conversely, community-associated factors (community engagement 0%, endorsement 0%, source reputation 0%) are not valued by users with less than a year of experience, but gain attention with increased experience, though still less than model-related attributes. This suggests that users are more dependent on direct model interaction than on external validation.
A clear trend also emerges based on user experience: experienced users place less emphasis on the model’s controllability (6%) and ethical compliance . This could imply that, with increasing familiarity, users developed a level of trust that mitigates concerns about how to use the tool effectively and the model’s benevolence. However, experienced users increasingly value the interpretability (56%) of the generated output and prefer models that integrate seamlessly into their established procedure.
Figure 5. Trust Antecedents Selection Based on Years of Experience in SE: Trust Antecedents
We observe a similar result from Fig. 5, a heat map between users’ year of expertise in SE and their trust antecedent selection. However, we find that there is no significant relationship between higher levels of knowledge in SE and their selection of community-related factors. Interestingly, we observe that intermediates pay significantly less emphasis on robustness compared to novices and experts.
@@ 6.3. Complementary Analysis of Trust Antecedents
Trust in AI systems, including LLM, is shaped by multiple interrelated factors, including technical performance, user perception, and ethical considerations. In AI research, accuracy is widely recognized as a fundamental trust factor in all domains (Khandelwal et al., 2020; Hendrycks et al., 2021; Yin et al., 2023). In SE, ensuring accuracy and reliability is essential to prevent security vulnerabilities (Hou and Jansen, 2022; Ahuja et al., 2020; Hassan et al., 2024; Baldassarre et al., 2024a; Serban et al., 2021; Rutinowski et al., 2024). However, trust is not solely dependent on accuracy; overconfidence in the accuracy of the model can hinder trust by making unreliable results appear convincing (Huang et al., 2023). This phenomenon is observed in various AI applications, where users may develop misplaced trust in systems despite their inherent unreliability (Banovic et al., 2023).
@@
Beyond precision, trust is also influenced by the model’s ability to balance technical performance with inclusivity and multilingual capabilities. Natural Language Processing (NLP) research highlights that ensuring unbiased and culturally aware results remains a challenge when optimizing high performance in different linguistic contexts (Cabrero-Daniel and Sanagustín Cabrero, 2023; Zhang et al., 2023; Wang et al., 2024b). This trade-off between precision and inclusivity is particularly relevant for SE, where domain-specific terminology may not generalize well across different programming cultures.
@@
Robustness is another critical factor in trust formation. AI models, including LLMs, often struggle to maintain consistent performance on similar inputs. Studies show that even minor perturbations in prompts can lead to significantly different outputs (Ye et al., 2023; Mehrotra et al., 2024), and evaluations such as PromptBench highlight the sensitivity of LLMs to word-level variations (Zhu et al., 2024). Similar robustness challenges have been identified in NLP tasks, where adversarial manipulations can undermine the reliability of the model (Wang et al., 2022; Chen and Hsieh, 2023).
@@
Transparency and accountability further shape trust in AI systems. Research suggests that users develop greater trust when they perceive AI as competent, benevolent, and transparent in their decision-making (Baughan et al., 2023; Jakesch et al., 2019; Zerick et al., 2024; Gambo et al., 2024). However, the black-box nature of LLMs and corporate secrecy hinder transparency (Wu et al., 2022; Liao and Wortman Vaughan, 2024; Langer et al., 2021). Techniques such as uncertainty expressions (e.g.,’I am not sure’) can improve trust by signaling limitations (Kim et al., 2024), while XAI approaches offer interpretability at the potential cost of performance (Wu et al., 2022). The regulatory landscape also plays a role, with frameworks such as the EU AI Act emphasizing responsibility (EUw, [n. d.]; Jobin et al., 2019; Floridi and Cowls, 2019; Nissenbaum, 1996; Stettinger et al., 2024), yet leaving questions of liability for AI errors unresolved (Nissenbaum, 1996; Lassiter and Fleischmann, 2024). The proposed solutions, such as blockchain-based traceability, could improve accountability in AI decision making (Zhang et al., 2024).
@@
Finally, ethical considerations remain an essential, albeit sometimes secondary, factor in trust formation. Although performance and usability often take priority, fairness in AI is critical to long-term trustworthiness. Studies reveal that LLMs can introduce social biases even in code generation tasks (Liu et al., 2023), and research on models such as Codex indicates a trade-off between fairness and performance (Cui et al., 2024). Trust is also shaped by concerns about the impact of artificial intelligence on professional development. Similarly to educational AI, where trust is linked to fairness and correctness in recommendations (Amoozadeh et al., 2024), SE users worry that overreliance on LLM could degrade their programming skills. However, while not a primary focus for experienced users, ethical concerns remain an important aspect of long-term trust in AI (Baldassarre et al., 2024b).